# uniINF: Best-of-Both-Worlds Algorithm for Parameter-Free Heavy-Tailed MABs

### v1

@@ 1 Introduction
Multi-Armed Bandits (MAB) problem serves as a solid theoretical formulation for addressing the exploration-exploitation trade-off inherent in online learning. Existing research in this area often assumes sub-Gaussian loss (or reward) distributions (Lattimore and Szepesvári,, 2020) or even bounded ones (Auer et al.,, 2002). However, recent empirical evidences revealed that Heavy-Tailed (HT) distributions appear frequently in realistic tasks such as network routing (Liebeherr et al.,, 2012), algorithm portfolio selection (Gagliolo and Schmidhuber,, 2011), and online deep learning (Zhang et al.,, 2020). Such observations underscore the importance of developing MAB solutions that are robust to heavy-tailed distributions.
In this paper, we consider the Heavy-Tailed Multi-Armed Bandits (HTMAB) proposed by Bubeck et al., (2013). In this scenario, the loss distributions associated with each arm do not allow bounded variances but instead have their $\alpha$-th moment bounded by some constant $\sigma^{\alpha}$, where $\alpha\in(1,2]$ and $\sigma\geq 0$ are predetermined constants. Mathematically, we assume $\mathbb{E}[\lvert\ell_{i}\rvert^{\alpha}]\leq\sigma^{\alpha}$ for every arm $i$. Although numerous existing HTMAB algorithms operated under the assumption that the parameters $\sigma$ and $\alpha$ are known (Bubeck et al.,, 2013; Yu et al.,, 2018; Wei and Srivastava,, 2020), real-world applications often present limited knowledge of the true environmental parameters. Thus we propose to explore the scenario where the algorithm lacks any prior information about the parameters. This setup, a variant of classical HTMABs, is referred to as the Parameter-Free HTMAB problem.
@@
When the environments can be potentially adversarial, a desirable property for bandit algorithms is called Best-of-Both-Worlds (BoBW), as proposed by Bubeck and Slivkins, (2012). A BoBW algorithm behaves well in both stochastic and adversarial environments without knowing whether stochasticity is satisfied. More specifically, while ensuring near-optimal regret in adversarial environments, the algorithmic performance in a stochastic environment should automatically be boosted, ideally matching the optimal performance of those algorithms specially crafted for stochastic environments.
In practical applications of machine learning and decision-making, acquiring prior knowledge about the environment is often a considerable challenge (Talaei Khoei and Kaabouch,, 2023). This is not only regarding the distribution properties of rewards or losses, which may not conform to idealized assumptions such as sub-Gaussian or bounded behaviors, but also about the stochastic or adversarial settings the agent palying in. Such environments necessitate robust solutions that can adapt without prior distributional knowledge. Therefore, the development of a BoBW algorithm that operates effectively without this prior information – termed a parameter-free HTMAB algorithm – is not just a theoretical interest but a practical necessity.
Various previous work has made strides in enhancing the robustness and applicability of HTMAB algorithms. For instance, Huang et al., (2022) pioneered the development of the first BoBW algorithm for the HTMAB problem, albeit requiring prior knowledge of the heavy-tail parameters $(\alpha,\sigma)$ to achieve near-optimal regret guarantees. Genalti et al., (2024) proposed an Upper Confidence Bound (UCB) based parameter-free HTMAB algorithm, specifically designed for stochastic environments. Both algorithms’ regret adheres to the instance-dependent and instance-independent lower bounds established by Bubeck et al., (2013). Though excited progress has been made, the following important question still stays open, which further eliminates the need of prior knowledge about the environment:
@@
Our contribution. In this paper, we answer this open question affirmatively by designing a single algorithm uniINF that enjoys both Parameter-Free and BoBW properties — that is, it i) does not require any prior knowledge of the environment, e.g., $\alpha$ or $\sigma$, and ii) its performance when deployed in an adversarial environment nearly matches the universal instance-independent lower bound given by Bubeck et al., (2013), and it attains the instance-dependent lower bound in stochastic environments as well. For more details, we summarize the advantages of our algorithm in Table 1. Our research directly contributes to enhancing the robustness and applicability of bandit algorithms in a variety of unpredictable and non-ideal conditions. Our main contributions are three-fold:
@@
Lower Bound [3pt](Bubeck et al.,, 2013) $\Omega\left(\sum_{i\neq i^{\ast}}(\frac{\sigma^{\alpha}}{\Delta_{i}})^{\frac{1}{\alpha-1}}\log T\right)$
@@
RobustUCB [3pt](Bubeck et al.,, 2013) Only Stoc. $\operatorname{\mathcal{O}}\left(\sum_{i\neq i^{\ast}}(\frac{\sigma^{\alpha}}{\Delta_{i}})^{\frac{1}{\alpha-1}}\log T\right)$
@@
Robust MOSS [3pt](Wei and Srivastava,, 2020) Only Stoc. $\operatorname{\mathcal{O}}\left(\sum_{i\neq i^{\ast}}(\frac{\sigma^{\alpha}}{\Delta_{i}})^{\frac{1}{\alpha-1}}\log(\frac{T}{K}(\frac{\sigma^{\alpha}}{\Delta_{i}^{\alpha}})^{\frac{1}{\alpha-1}})\right)$ 22footnotemark: 2  [^3] Up to $\log(\sigma^{\alpha})$ and $\log(1/\Delta_{i}^{\alpha})$ factors.
@@
APE2 [3pt] (Lee et al.,, 2020) Only Stoc. ${\mathcal{O}}\left(e^{\sigma}+\sum_{i\neq i^{\ast}}(\frac{1}{\Delta_{i}})^{\frac{1}{\alpha-1}}(T\Delta_{i}^{\frac{\alpha}{\alpha-1}}\log K)^{\frac{\alpha}{(\alpha-1)\log K}}\right)$
@@
HTINF [3pt](Huang et al.,, 2022) Stoc. $\operatorname{\mathcal{O}}\left(\sum_{i\neq i^{\ast}}(\frac{\sigma^{\alpha}}{\Delta_{i}})^{\frac{1}{\alpha-1}}\log T\right)$
@@
OptHTINF [3pt](Huang et al.,, 2022) Stoc. $\operatorname{\mathcal{O}}\left(\sum_{i\neq i^{\ast}}(\frac{\sigma^{2\alpha}}{\Delta_{i}^{3-\alpha}})^{\frac{1}{\alpha-1}}\log T\right)$
@@
AdaTINF [3pt] (Huang et al.,, 2022) Only Adv. $\operatorname{\mathcal{O}}\left(\sigma K^{1-\nicefrac{{1}}{{\alpha}}}T^{\nicefrac{{1}}{{\alpha}}}\right)$
AdaR-UCB [3pt](Genalti et al.,, 2024) Only Stoc. $\operatorname{\mathcal{O}}\left(\sum_{i\neq i^{\ast}}(\frac{\sigma^{\alpha}}{\Delta_{i}})^{\frac{1}{\alpha-1}}\log T\right)$
@@ 2 Related Work
Heavy-Tailed Multi-Armed Bandits.
HTMABs were introduced by Bubeck et al., (2013), who gave both instance-dependent and instance-independent lower and upper bounds under stochastic assumptions. Various efforts have been devoted in this area since then. To exemplify, Wei and Srivastava, (2020) removed a sub-optimal $(\log T)^{1-1/\alpha}$ factor in the instance-independent upper bound; Yu et al., (2018) developed a pure exploration algorithm for HTMABs; Medina and Yang, (2016), Kang and Kim, (2023), and Xue et al., (2024) considered the linear HTMAB problem; and Dorn et al., (2024) specifically investigated the case where the heavy-tailed reward distributions are presumed to be symmetric. Nevertheless, all these works focused on stochastic environments and required the prior knowledge of heavy-tail parameters $(\alpha,\sigma)$. In contrast, this paper focuses on both-of-both-worlds algorithms in parameter-free HTMABs, which means i) the loss distributions can possibly be non-stationary, and ii) the true heavy-tail parameters $(\alpha,\sigma)$ remain unknown.
Best-of-Both-Worlds Algorithms.
Bubeck and Slivkins, (2012) pioneered the study of Best-of-Both-Worlds bandit algorithms, and was followed by various improvements including EXP3-based approaches (Seldin and Slivkins,, 2014; Seldin and Lugosi,, 2017), FTPL-based approaches (Honda et al.,, 2023; Lee et al.,, 2024), and FTRL-based approaches (Wei and Luo,, 2018; Zimmert and Seldin,, 2019; Jin et al.,, 2023). When the loss distribution can be heavy-tailed, Huang et al., (2022) gave an algorithm HTINF that achieves best-of-both-worlds property under the known-$(\alpha,\sigma)$ assumption. Unfortunately, without access to these true parameters, their alternative algorithm OptHTINF failed to achieve near-optimal regret guarantees in either adversarial or stochastic environments.
Parameter-Free HTMABs.
Another line of research aimed at getting rid of the prior knowledge of $\alpha$ or $\sigma$, which we call Parameter-Free HTMABs. Along this line, Kagrecha et al., (2019) presented the GSR method to identify the optimal arm in HTMAB without any prior knowledge. In terms of regret minimization, Lee et al., (2020) and Lee and Lim, (2022) considered the case when $\sigma$ is unknown. Genalti et al., (2024) were the first to achieve the parameter-free property while maintaining near-optimal regret. However, all these algorithms fail in adversarial environments – not to mention the best-of-both-worlds property which requires optimality in both stochastic and adversarial environments.
@@ 3 Preliminaries: Heavy-Tailed Multi-Armed Bandits
Notations.
For an integer $n\geq 1$, $[n]$ denotes the set $\{1,2,\ldots,n\}$. For a finite set $\mathcal{X}$, $\Delta(\mathcal{X})$ denotes the set of probability distributions over $\mathcal{X}$, often also referred to as the simplex over $\mathcal{X}$. We also use $\Delta^{[K]}:=\Delta([K])$ to denote the simplex over $[K]$. We use $\operatorname{\mathcal{O}}$ to hide all constant factors, and use $\operatorname{\widetilde{\mathcal{O}}}$ to additionally suppress all logarithmic factors. We usually use bold letters ${\bm{x}}$ to denote a vector, while $x_{i}$ denotes an entry of the vector. Unless mentioned explicitly, $\log(x)$ denotes the natural logarithm of $x$. Throughout the text, we will use $\{\mathcal{F}_{t}\}_{t=0}^{T}$ to denote the natural filtration, i.e., $\mathcal{F}_{t}$ represents the $\sigma$-algebra generated by all random observations made during the first $t$ time-slots.
@@
In Heavy-Tailed MABs (HTMAB), for every $t\in[T]$ and $i\in[K]$, the loss is independently sampled from some heavy-tailed distribution $\nu_{t,i}$ in the sense that $\mathbb{E}_{\ell\sim\nu_{t,i}}[\lvert\ell\rvert^{\alpha}]\leq\sigma^{\alpha}$, where $\alpha\in(1,2]$ and $\sigma\geq 0$ are some pre-determined but unknown constants. A Best-of-Both-Worlds (BoBW) algorithm is one that behaves well in both stochastic and adversarial environments (Bubeck and Slivkins,, 2012), where stochastic environments are those with time-homogeneous $\{\bm{\nu}_{t}\}_{t\in[T]}$ (i.e., $\nu_{t,i}=\nu_{1,i}$ for every $t\in[T]$ and $i\in[K]$) and adversarial environments are those where $\nu_{t,i}$’s can depend on both $t$ and $i$. However, we do not allow the loss distributions to depend on the player’s previous actions. [^6] Called oblivious adversary model (Bubeck and Slivkins,, 2012; Wei and Luo,, 2018; Zimmert and Seldin,, 2019).
Before concluding this section, we make the following essential assumption. As shown in (Genalti et al.,, 2024, Theorems 2 & 3), without Assumption 1, there does not exist HTMAB algorithms that can match the worst-case regret guarantees (Bubeck et al.,, 2013) without knowing either $\alpha$ or $\sigma$.
Assumption 1 (Truncated Non-Negative Loss (Huang et al.,, 2022, Assumption 3.6)).
@@
Additionally, we make the following assumption for stochastic cases. Assumption 2 is common for algorithms utilizing self-bounding analyses, especially those with BoBW properties (Gaillard et al.,, 2014; Luo and Schapire,, 2015; Wei and Luo,, 2018; Zimmert and Seldin,, 2019; Ito et al.,, 2022).
@@ 4 The BoBW HTMAB Algorithm uniINF
7:     Calculate the importance sampling estimate of ${\bm{\ell}}^{\text{skip}}_{t}$, namely ~$\widetilde{{\bm{\ell}}}_{t}$, where
~$\widetilde{\ell}_{t,i}=\frac{\ell^{\text{skip}}_{t,i}}{x_{t,i}}\cdot\mathbbm{1}[i=i_{t}],\quad\forall i\in[K].$
@@ 4.1 Refined Log-Barrier Analysis
We adopt the log-barrier regularizer $\Psi_{t}({\bm{x}}):=-S_{t}\sum_{i=1}^{K}\log x_{i}$ in Equation 2 where $S_{t}^{-1}$ is the learning rate in round $t$. While log-barrier regularizers were commonly used in the literature for data-adaptive bounds such as small-loss bounds (Foster et al.,, 2016), path-length bounds (Wei and Luo,, 2018), and second-order bounds (Ito,, 2021), this paper introduces novel analysis illustrating that log-barrier regularizers also provide environment-adaptivity for both stochastic and adversarial settings.
Precisely, it is known that log-barrier applied to a loss sequence $\{\bm{c}_{t}\}_{t=1}^{T}$ ensures $\sum_{t=1}^{T}\langle\bm{x}_{t}-\bm{y},\bm{c}_{t}\rangle\lesssim\sum_{t=1}^{T}((S_{t+1}-S_{t})K\log T+\textsc{Div}_{t})$ where $\textsc{Div}_{t}\leq S_{t}^{-1}\sum_{i=1}^{K}{\color[rgb]{0,0.5,0}x_{t,i}^{2}}c_{t,i}^{2}$ for non-negative $\bm{c}_{t}$’s (Foster et al.,, 2016, Lemma 16) and $\textsc{Div}_{t}\leq S_{t}^{-1}\sum_{i=1}^{K}{\color[rgb]{0,0.5,0}x_{t,i}}c_{t,i}^{2}$ for general $\bm{c}_{t}$’s (Dai et al.,, 2023, Lemma 3.1). In comparison, our Lemmas 4 and 5 focus on the case where $S_{t}$ is adequately large compared to $\lVert\bm{c}_{t}\rVert_{\infty}$ and give a refined version of $\textsc{Div}_{t}\leq S_{t}^{-1}\sum_{i=1}^{K}{\color[rgb]{0,0.5,0}x_{t,i}^{2}(1-x_{t,i})^{2}}c_{t,i}^{2}$, which means
@@
The extra $(1-x_{t,i})^{2}$ terms are essential to exclude the optimal arm $i^{\ast}\in[K]$ — a nice property that leads to best-of-both-worlds guarantees (Zimmert and Seldin,, 2019; Dann et al., 2023a, ). To give more technical details on why we need this $(1-x_{t,i})^{2}$, in Section 5.1, we will decompose the regret into skipping error $\sum_{t=1}^{T}(\ell_{t,i_{t}}-\ell_{t,i_{t}}^{{\text{skip}}})\mathbbm{1}[i_{t}\neq i^{\ast}]$ and main regret which is roughly $\sum_{t=1}^{T}\langle\bm{x}_{t}-\bm{y},\bm{\ell}^{{\text{skip}}}_{t}\rangle$. The skipping errors already include the indicator $\mathbbm{1}[i_{t}\neq i^{\ast}]$, so the exclusion of $i^{\ast}$ is automatic. However, for the main regret, we must manually introduce some $(1-x_{t,i})$ to exclude $i^{\ast}$ — which means the previous bounds mentioned above do not apply, while our novel Equation 5 is instead helpful.

### v2

@@ 1 Introduction
Multi-Armed Bandits (MAB) problem serves as a solid theoretical formulation for addressing the exploration-exploitation trade-off inherent in online learning. Existing research in this area often assumes sub-Gaussian loss (or reward) distributions (Lattimore & Szepesvári, 2020) or even bounded ones (Auer et al., 2002). However, recent empirical evidences revealed that Heavy-Tailed (HT) distributions appear frequently in realistic tasks such as network routing (Liebeherr et al., 2012), algorithm portfolio selection (Gagliolo & Schmidhuber, 2011), and online deep learning (Zhang et al., 2020). Such observations underscore the importance of developing MAB solutions that are robust to heavy-tailed distributions.
In this paper, we consider the Heavy-Tailed Multi-Armed Bandits (HTMAB) proposed by Bubeck et al. (2013). In this scenario, the loss distributions associated with each arm do not allow bounded variances but instead have their $\alpha$-th moment bounded by some constant $\sigma^{\alpha}$, where $\alpha\in(1,2]$ and $\sigma\geq 0$ are predetermined constants. Mathematically, we assume $\mathbb{E}[\lvert\ell_{i}\rvert^{\alpha}]\leq\sigma^{\alpha}$ for every arm $i$. Although numerous existing HTMAB algorithms operated under the assumption that the parameters $\sigma$ and $\alpha$ are known (Bubeck et al., 2013; Yu et al., 2018; Wei & Srivastava, 2020), real-world applications often present limited knowledge of the true environmental parameters. Thus we propose to explore the scenario where the algorithm lacks any prior information about the parameters. This setup, a variant of classical HTMABs, is referred to as the Parameter-Free HTMAB problem.
@@
When the environments can be potentially adversarial, a desirable property for bandit algorithms is called Best-of-Both-Worlds (BoBW), as proposed by Bubeck & Slivkins (2012). A BoBW algorithm behaves well in both stochastic and adversarial environments without knowing whether stochasticity is satisfied. More specifically, while ensuring near-optimal regret in adversarial environments, the algorithmic performance in a stochastic environment should automatically be boosted, ideally matching the optimal performance of those algorithms specially crafted for stochastic environments.
In practical applications of machine learning and decision-making, acquiring prior knowledge about the environment is often a considerable challenge (Talaei Khoei & Kaabouch, 2023). This is not only regarding the distribution properties of rewards or losses, which may not conform to idealized assumptions such as sub-Gaussian or bounded behaviors, but also about the stochastic or adversarial settings the agent palying in. Such environments necessitate robust solutions that can adapt without prior distributional knowledge. Therefore, the development of a BoBW algorithm that operates effectively without this prior information – termed a parameter-free HTMAB algorithm – is not just a theoretical interest but a practical necessity.
Various previous work has made strides in enhancing the robustness and applicability of HTMAB algorithms. For instance, Huang et al. (2022) pioneered the development of the first BoBW algorithm for the HTMAB problem, albeit requiring prior knowledge of the heavy-tail parameters $(\alpha,\sigma)$ to achieve near-optimal regret guarantees. Genalti et al. (2024) proposed an Upper Confidence Bound (UCB) based parameter-free HTMAB algorithm, specifically designed for stochastic environments. Both algorithms’ regret adheres to the instance-dependent and instance-independent lower bounds established by Bubeck et al. (2013). Though excited progress has been made, the following important question still stays open, which further eliminates the need of prior knowledge about the environment:
@@
Our contribution. In this paper, we answer this open question affirmatively by designing a single algorithm uniINF that enjoys both Parameter-Free and BoBW properties — that is, it i) does not require any prior knowledge of the environment, e.g., $\alpha$ or $\sigma$, and ii) its performance when deployed in an adversarial environment nearly matches the universal instance-independent lower bound given by Bubeck et al. (2013), and it attains the instance-dependent lower bound in stochastic environments as well. For more details, we summarize the advantages of our algorithm in Table 1. Our research directly contributes to enhancing the robustness and applicability of bandit algorithms in a variety of unpredictable and non-ideal conditions. Our main contributions are three-fold:
@@
Lower Bound [3pt](Bubeck et al., 2013) $\Omega\left(\sum_{i\neq i^{\ast}}(\frac{\sigma^{\alpha}}{\Delta_{i}})^{\frac{1}{\alpha-1}}\log T\right)$
@@
RobustUCB [3pt](Bubeck et al., 2013) Only Stoc. $\operatorname{\mathcal{O}}\left(\sum_{i\neq i^{\ast}}(\frac{\sigma^{\alpha}}{\Delta_{i}})^{\frac{1}{\alpha-1}}\log T\right)$
@@
Robust MOSS [3pt](Wei & Srivastava, 2020) Only Stoc. $\operatorname{\mathcal{O}}\left(\sum_{i\neq i^{\ast}}(\frac{\sigma^{\alpha}}{\Delta_{i}})^{\frac{1}{\alpha-1}}\log(\frac{T}{K}(\frac{\sigma^{\alpha}}{\Delta_{i}^{\alpha}})^{\frac{1}{\alpha-1}})\right)$ 22footnotemark: 2  [^3] Up to $\log(\sigma^{\alpha})$ and $\log(1/\Delta_{i}^{\alpha})$ factors.
@@
APE2 [3pt] (Lee et al., 2020b) Only Stoc. ${\mathcal{O}}\left(e^{\sigma}+\sum_{i\neq i^{\ast}}(\frac{1}{\Delta_{i}})^{\frac{1}{\alpha-1}}(T\Delta_{i}^{\frac{\alpha}{\alpha-1}}\log K)^{\frac{\alpha}{(\alpha-1)\log K}}\right)$
@@
HTINF [3pt](Huang et al., 2022) Stoc. $\operatorname{\mathcal{O}}\left(\sum_{i\neq i^{\ast}}(\frac{\sigma^{\alpha}}{\Delta_{i}})^{\frac{1}{\alpha-1}}\log T\right)$
@@
OptHTINF [3pt](Huang et al., 2022) Stoc. $\operatorname{\mathcal{O}}\left(\sum_{i\neq i^{\ast}}(\frac{\sigma^{2\alpha}}{\Delta_{i}^{3-\alpha}})^{\frac{1}{\alpha-1}}\log T\right)$
@@
AdaTINF [3pt] (Huang et al., 2022) Only Adv. $\operatorname{\mathcal{O}}\left(\sigma K^{1-\nicefrac{{1}}{{\alpha}}}T^{\nicefrac{{1}}{{\alpha}}}\right)$
AdaR-UCB [3pt](Genalti et al., 2024) Only Stoc. $\operatorname{\mathcal{O}}\left(\sum_{i\neq i^{\ast}}(\frac{\sigma^{\alpha}}{\Delta_{i}})^{\frac{1}{\alpha-1}}\log T\right)$
@@ 2 Related Work
Due to space limitations, we defer the detailed comparison with the related works into Appendix A and only discuss a few most related ones here.
Heavy-Tailed Multi-Armed Bandits. HTMABs were introduced by Bubeck et al. (2013), who gave both instance-dependent and independent lower and upper bounds under stochastic assumptions. Various efforts have been devoted in this area since then. To exemplify, Wei & Srivastava (2020) improve the instance-independent upper bound; Yu et al. (2018) developed a pure exploration algorithm for HTMABs; Medina & Yang (2016); Kang & Kim (2023); Xue et al. (2024) considered the linear HTMAB problem; and Dorn et al. (2024) investigated the symmetric reward distribution.
Best-of-Both-Worlds Algorithms. Bubeck & Slivkins (2012) pioneered the study of BoBW bandit algorithms, and was followed by various improvements including EXP3-based approaches (Seldin & Slivkins, 2014), FTPL-based approaches (Honda et al., 2023), and FTRL-based approaches (Wei & Luo, 2018; Zimmert & Seldin, 2019). When the loss distribution can be heavy-tailed, Huang et al. (2022) achieves best-of-both-worlds property under the known-$(\alpha,\sigma)$ assumption.
Parameter-Free HTMABs. Another line of HTMABs’ research aimed at getting rid of the prior knowledge of $\alpha$ or $\sigma$, which we call Parameter-Free HTMABs. Along this line, Kagrecha et al. (2019) presented the GSR method to identify the optimal arm in HTMAB without any prior knowledge. In terms of regret minimization, Lee et al. (2020b) and Lee & Lim (2022) considered the case when $\sigma$ is unknown. Genalti et al. (2024) were the first to achieve the parameter-free property while maintaining near-optimal regret. However, all these algorithms fail in adversarial environments.
@@ 3 Preliminaries: Heavy-Tailed Multi-Armed Bandits
Notations. For an integer $n\geq 1$, $[n]$ denotes the set $\{1,2,\ldots,n\}$. For a finite set $\mathcal{X}$, $\Delta(\mathcal{X})$ denotes the set of probability distributions over $\mathcal{X}$, often also referred to as the simplex over $\mathcal{X}$. We also use $\Delta^{[K]}:=\Delta([K])$ to denote the simplex over $[K]$. We use $\operatorname{\mathcal{O}}$ to hide all constant factors, and use $\operatorname{\widetilde{\mathcal{O}}}$ to additionally suppress all logarithmic factors. We usually use bold letters ${\bm{x}}$ to denote a vector, while $x_{i}$ denotes an entry of the vector. Unless mentioned explicitly, $\log(x)$ denotes the natural logarithm of $x$. Throughout the text, we will use $\{\mathcal{F}_{t}\}_{t=0}^{T}$ to denote the natural filtration, i.e., $\mathcal{F}_{t}$ represents the $\sigma$-algebra generated by all random observations made during the first $t$ time-slots.
@@
In Heavy-Tailed MABs (HTMAB), for every $t\in[T]$ and $i\in[K]$, the loss is independently sampled from some heavy-tailed distribution $\nu_{t,i}$ in the sense that $\mathbb{E}_{\ell\sim\nu_{t,i}}[\lvert\ell\rvert^{\alpha}]\leq\sigma^{\alpha}$, where $\alpha\in(1,2]$ and $\sigma\geq 0$ are some pre-determined but unknown constants. A Best-of-Both-Worlds (BoBW) algorithm is one that behaves well in both stochastic and adversarial environments (Bubeck & Slivkins, 2012), where stochastic environments are those with time-homogeneous $\{\bm{\nu}_{t}\}_{t\in[T]}$ (i.e., $\nu_{t,i}=\nu_{1,i}$ for every $t\in[T]$ and $i\in[K]$) and adversarial environments are those where $\nu_{t,i}$’s can depend on both $t$ and $i$. However, we do not allow the loss distributions to depend on the player’s previous actions. [^6] Called oblivious adversary model (Bubeck & Slivkins, 2012; Wei & Luo, 2018; Zimmert & Seldin, 2019).
Before concluding this section, we make the following important assumption which is also considered in Huang et al. (2022) and Genalti et al. (2024). Notice that Assumption 1 is strictly weaker than the common "non-negative losses" assumption in the MABs literature (Auer et al., 2002; Lee et al., 2020a; Jin et al., 2023), which only needs the truncated non-negativity for the optimal arm. We make a detailed discussion on this assumption in Appendix B.
Assumption 1 (Truncated Non-Negative Loss (Huang et al., 2022, Assumption 3.6)).
@@
Additionally, we make the following assumption for stochastic cases. Assumption 2 is common for algorithms utilizing self-bounding analyses, especially those with BoBW properties (Gaillard et al., 2014; Luo & Schapire, 2015; Wei & Luo, 2018; Zimmert & Seldin, 2019; Ito et al., 2022). [^7] While Assumption 2 is commonly adopted in prior work, it is worth noting that some Tsallis-INF algorithms do not require this assumption as analyzed in Ito (2021a) and further extended in Jin et al. (2023).
@@ 4 The BoBW HTMAB Algorithm uniINF
7:     Calculate the importance sampling estimate of ${\bm{\ell}}^{\text{skip}}_{t}$, namely ~$\widetilde{{\bm{\ell}}}_{t}$, where ~$\widetilde{\ell}_{t,i}=\frac{\ell^{\text{skip}}_{t,i}}{x_{t,i}}\cdot\mathbbm{1}[i=i_{t}],\quad\forall i\in[K].$
@@ 4.1 Refined Log-Barrier Analysis
We adopt the log-barrier regularizer $\Psi_{t}({\bm{x}}):=-S_{t}\sum_{i=1}^{K}\log x_{i}$ in Equation 2 where $S_{t}^{-1}$ is the learning rate in round $t$. While log-barrier regularizers were commonly used in the literature for data-adaptive bounds such as small-loss bounds (Foster et al., 2016), path-length bounds (Wei & Luo, 2018), and second-order bounds (Ito, 2021b), and have also been shown to enable the best-of-both-worlds property in certain settings, this paper introduces novel analysis illustrating that log-barrier regularizers also provide environment-adaptivity for both stochastic and adversarial settings.
Precisely, it is known that log-barrier applied to a loss sequence $\{\bm{c}_{t}\}_{t=1}^{T}$ ensures $\sum_{t=1}^{T}\langle\bm{x}_{t}-\bm{y},\bm{c}_{t}\rangle\lesssim\sum_{t=1}^{T}((S_{t+1}-S_{t})K\log T+\textsc{Div}_{t})$ where $\textsc{Div}_{t}\leq S_{t}^{-1}\sum_{i=1}^{K}{\color[rgb]{0,0.5,0}x_{t,i}^{2}}c_{t,i}^{2}$ for non-negative $\bm{c}_{t}$’s (Foster et al., 2016, Lemma 16) and $\textsc{Div}_{t}\leq S_{t}^{-1}\sum_{i=1}^{K}{\color[rgb]{0,0.5,0}x_{t,i}}c_{t,i}^{2}$ for general $\bm{c}_{t}$’s (Dai et al., 2023, Lemma 3.1). In comparison, our Lemmas 4 and 5 focus on the case where $S_{t}$ is adequately large compared to $\lVert\bm{c}_{t}\rVert_{\infty}$ and give a refined version of $\textsc{Div}_{t}\leq S_{t}^{-1}\sum_{i=1}^{K}{\color[rgb]{0,0.5,0}x_{t,i}^{2}(1-x_{t,i})^{2}}c_{t,i}^{2}$, which means
@@
The extra $(1-x_{t,i})^{2}$ terms are essential to exclude the optimal arm $i^{\ast}\in[K]$ — a nice property that leads to best-of-both-worlds guarantees (Zimmert & Seldin, 2019; Dann et al., 2023b). To give more technical details on why we need this $(1-x_{t,i})^{2}$, in Section 5.1, we will decompose the regret into skipping error $\sum_{t=1}^{T}(\ell_{t,i_{t}}-\ell_{t,i_{t}}^{{\text{skip}}})\mathbbm{1}[i_{t}\neq i^{\ast}]$ and main regret which is roughly $\sum_{t=1}^{T}\langle\bm{x}_{t}-\bm{y},\bm{\ell}^{{\text{skip}}}_{t}\rangle$. The skipping errors already include the indicator $\mathbbm{1}[i_{t}\neq i^{\ast}]$, so the exclusion of $i^{\ast}$ is automatic. However, for the main regret, we must manually introduce some $(1-x_{t,i})$ to exclude $i^{\ast}$ — which means the previous bounds mentioned above do not apply, while our novel Equation 5 is instead helpful.